OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

y = -2x + 19 y = x + 7 Just use substitution where you take a variable and plug in an equivalent. It plainly says that y is equal to -2x + 19, so use that for y in the second equation. -2x + 19 = x + 7 -3x = -12 x = 4 Then, plug this into the equation. y = x + 7 y = 4 + 7 y = 11

OpenStudy (anonymous):

Yup. So if a=b and a=c. Then b=c. That's the key. :D